# Design and Performance of a Universal SiPM Readout System for X- and Gamma-Ray Missions

### v1

@@ 2.2.5 ASICs
The two different amplifiers as well as the two thresholds allow to setup a complex trigger scheme. The system was originally designed for a gamma-ray polarimeter for which coincident events between at least 2 channels are of interest. These 2 detector channels can be within one 64-channel front-end readout, or can be between 2 different front-ends. To accommodate this, the time threshold and charge threshold are used for different trigger types. The charge trigger is setup at a higher level such that, in case of 1 channel having a significant energy deposition, the FPGA initiates a readout of the ADCs. In addition, the FPGA sends out a signal on the ’trig out’ line to the back-end thereby allowing the back-end to combine the produced trigger packets from this front-end with those from other front-ends if these have recorded a trigger within a set coincidence time window (set to 100 ns for POLAR-2). This allows to detect coincident events within the full detector, while setting the threshold well above the dark noise region avoids dark noise induced dead time.
@@
The time threshold is instead setup to record coincident events within a single front-end. The time threshold is set at level within the dark noise region (typically around 4 or 5 photo-electrons). However, to avoid significant dead time, the ADC readout is only initialized when at least 2 of the time trigger lines are high. This check is performed by the FPGA using the multiplexed time threshold outputs from the 2 Citirocs. This check is performed within 90 nanoseconds. The multiplicity counting by the FPGA also allows to reduce charged particle induced events, as these typically produce events within the detector with a high number of detector channels triggering. In case that within a single front-end the multiplicity is high, the FGPA uses one of the spare LVDS trigger lines to indicate this to the back-end, thereby allowing it to veto the event.
Finally, the ADC readout can be initialized by force from the back-end using the LVDS ’trig in’ line. This option can be used to measure the baseline for the 64 channels, as well as to measure the electronic noise levels in the detector.
@@
Apart from its capability to provide a complex, flexible trigger logic, the Citiroc 1A also allows to fine-tune a range of parameters for the 32 channels. This includes:
  • The high gain amplifier with 8 bits
  • The low gain amplifier with 8 bits
  • The charge threshold level with 6 bits (a common 32 channel charge threshold level is set with 10 bits)
  • The time threshold with 6 bits (a common 32 channel time threshold level is set with 10 bits)
  • The bias voltage with 10 bits on either a 2.5 or 4.5 V level.
Here, the bias voltage for the 32 channels is adapted by the ASIC by raising the voltage level on the input of the signal line. As such a common bias voltage can be applied to the 32 channels, while using the ASIC it can be adapted to accommodate variations in the breakdown between various SiPMs. As such variations, within a single type, are typically of the order of hundreds of mV, the 2.5 V option is normally sufficient to produce a uniform over-voltage. The 4.5 V option can be of use in case, for example, one wants to operate different channels at different over-voltage levels.
It should finally be noted that since the start of the development of this front-end, the Radioroc 2 ASIC has been developed by Weeroc (Contino et al., 2020). This ASIC has all the features of the Citiroc 1A while allowing it to read out 64 channels compared to the 32 from the Citiroc. An upgrade to this ASIC in future versions is therefore being considered.
@@ 2.2.6 HVPS
It should be noted that after radiation damage the dark current will significantly increase, see e.g. Mitchell et al. (2022); De Angelis et al. (2023); Merzi et al. (2024). For POLAR-2, the yearly radiation dose has been estimated to be 0.0789 Gy/yr (at an altitude of 383 km) through simulations. This would result in an increase in dark current of a factor of 7.2 (for the 6025 or 6050 types would be a factor 15.7 and 15.4, respectively (De Angelis et al., 2023)). After one year, the current in one readout system would therefore increase to the order of $1-10mA$, while after several years, when correcting for annealing, this could increase to the $100\,\mathrm{mA}$ level. The LT3482 DC/DC converter can provide up to 200 mA at 70 V over voltage. Therefore, it allows to continue to power the SiPM array even after significant levels of radiation damage.
@@ 3.2 Electronic Noise
Under normal conditions, the system would be triggered to perform a readout of all the ADC values when one, or potentially several, detector channels are above the charge or time threshold. However, in order to measure the baseline as well as the electronic noise, the system can be forced to trigger using the ’trig in’ LVDS line. During standard data taking this can be done, for example, once per second to monitor potential drifts in the baseline due to temperature. Here we use this function to measure the electronic noise. For this purpose approximately 5000 triggers were forced while the high voltage was switched off, thereby only measuring the baseline, and its variations for the 64 channels, both in the high gain and low gain output. An example of the ADC distribution of channel 0, fitted using a Gaussian function can be seen on the left in figure 4. The $\sigma$ values for the 64 channels, for the high gain, can be seen on the right. Similar values were found for the low gain, indicating a typical electronics-induced noise of the order of 5 ADC.
@@
The electronic noise can consist of two components, one which is shared between all the channels in an ASIC, or within the full board, and another which is individual for each channel. The presence of any shared noise component was tested by taking the mean ADC value of the 64 channels and studying its distribution. The resulting distribution was found to have a width consistent with 0 ADC, thereby excluding the presence of a significant shared noise component.
@@ 3.3 Rate Capability
The SiPM readout is originally designed to observe transient sources like GRBs. As such it is designed to handle relatively high photon rates. From the POLAR mission, the brightest observed GRB was the short GRB 170127C which, after correcting for dead time, induced a gamma-ray rate of  $\approx\,200\,\mathrm{ph\,cm^{-2}\,s^{-1}}$ for a duration of 100 ms. For the brightest GRB ever observed, 221009A the GRBAlpha collaboration found a flux of $1020\,\mathrm{ph\,cm^{-2}\,s^{-1}}$ in the 50-300 keV range (Ripa et al., 2023). In order for the detector to be capable of observing such transients without a significant amount of dead time, it should therefore be capable of handling trigger rates of hundreds of $\mathrm{triggers/(cm^{2}\,s)}$ assuming a few tens of percent triggering efficiency.. Assuming the baseline design, which reads out 64 SiPMs with dimensions of $6\times 6\,\mathrm{mm^{2}}$, this corresponds to being able to handle a trigger rate of the front-end board of the order of several kHz.
The main sources which limit the readout rate of the system are identified as:
  • The ADC readout time
  • The event packaging time in the FPGA
  • The readout rate of the board by the back-end
The 14-bit ADC readout is the first bottleneck and puts a lower limit on the possible time between two triggered events. With a clock speed of 1.5 MHz, the total readout time is $9.3\mu s$.
@@
The effects can be seen on the left in figure 5 which shows the time between two consecutive events for a data run where thresholds were set such that the system was triggering at high rates on dark noise from the SiPMs. A large times, the peak induced by the back-end at 2 ms can be observed, while at lower times the effects due to the readout system itself. These can be more clearly observed in the inset. The lower limit at $9\mu s$ can clearly be observed, while a kink at around $200\mu s$, induced by the FGPA can be observed. Fitting the region below the kink using an exponential we can derive a trigger rate of 4 kHz.
@@
To further verify the trigger rate, charges were directly injected into the setup using the charge injection point on the PCB. The charges, produced by a pulse generator, are injected directly into the ASICs, thereby allowing to test the full electronics chain. When injecting charges at a rate of 1–2 kHz, the correct time between triggers could be recovered. Only when exceeding 5 kHz do we start to see this being distorted. The result of injecting charges at 10 kHz can be observed on the right of figure 5. While for an ideal detector, we would expect to measure 100 $\mu s$ delays, here, for about $60\%$ of the cases, a delay of 200 $\mu s$ is measured, indicating that packets are being lost.
@@ 3.4 Crosstalk
A typical detector using scintillators, SiPMs, and electronics can suffer from various types of crosstalk. Firstly, there can be optical crosstalk between the scintillators. Secondly, crosstalk between microcells in the SiPM occurs. These first two forms of crosstalk depend on the detector design. The third, the electronic crosstalk, depends on the readout design. As the system discussed here is developed to read out 64 independent SiPMs, it is important to avoid any electronic crosstalk between readout lines in the system. To verify that no significant crosstalk of this type exists, data with large signals is required. For this purpose, we use data from acquisitions where the system was coupled to an array of GAGG crystals which was illuminated by a 100 keV beam. The correlation between ADC values in different channels was evaluated and studied for crosstalk. To verify that the analysis works, we initially looked at data from two SiPM channels which are geometrically adjacent. The mean ADC value as registered in a geometrically neighboring channel, as a function of the ADC value in the triggering channel, which here was illuminated by the beam, can be seen on the left in figure 6. We see a clear correlation between the two ADC values, indicating that a significant level of crosstalk, either optical or between SiPM channels, exists. This was expected, since the GAGG crystal array is not well optically insulated, thereby resulting in optical crosstalk of $\approx 10\%$.
@@
4 Performance with EJ-248M and MPPC S13661
The SiPM readout system was originally designed as part of the POLAR-2 mission, a large-scale plastic scintillator-based polarimeter (Kole et al., 2024). It was therefore initially optimized to read out a densely packed array of plastic scintillator bars. Although originally, EJ-200 was chosen for this purpose, it was found that EJ-248M performed better in terms of light yield despite its lower intrinsic scintillation light yield compared to EJ-200. This was found to be a result of EJ-248M being a relatively harder plastic and therefore allowing a smoother surface quality, resulting in a larger fraction of the light reaching the SiPM. The details of this are discussed in De Angelis et al. (2024), where the details on the optical coupling are also presented along with the calculation of the achieved light yield. In an optimized setup an average of $\approx 1.6$ p.e./keV was measured.
Similarly, the combination of the readout system together with an EJ-248M, as well as an EJ-200 array, was tested during a beam test at the European Synchrotron Radiation Facility (ESRF) of which the details are presented in Kole et al. (2024). While the details on the light yield and spectra can be found in Kole et al. (2024), it should be noted that those results were produced using the same electronics, though with a slightly faulty firmware. As a result, for the data presented there, peak sensing was not initiated for events where the charge trigger was not exceeded. As a result, the ADC values for non-triggering events were around the baseline level. Here, we present the typical response of the system after the firmware correction. For details on the performance of the system with a polarized beam, we refer the reader to Kole et al. (2024), while for details on the optical performance for this system, we refer the reader to De Angelis et al. (2024).
@@ 4.1 Spectral Measurements
In order to study the performance of the readout system with an array of EJ-248M scintillators, a full detector, consisting of 64 scintillator bars, was scanned using a collimated ${}^{241}Am$ source. The emission spectrum of this source is dominated by a $59.5\,\mathrm{keV}$ line. The spectrum from one of the detector channels can be seen in figure 7. In this spectrum, which is shown after pedestal subtraction was performed, events that have induced a charge trigger are shown in red, while those that did not are shown in green. It can be seen that the threshold was set so low that a small fraction of events from the dark noise region induce a trigger. This is possible as the threshold has a finite width resulting in a non-zero threshold efficiency at low ADC values. This, in combination with the very high rates in the dark noise region, results in such triggers. Increasing the threshold by approximately 50 ADC fully removes the dark noise-induced triggers.
Figure 7: The energy spectra, in ADC, after pedestal subtraction both in the high gain range (top) and low gain range (bottom). These spectra were obtained when irradiating the readout system connected to a plastic scintillator array at room temperature with an ${}^{241}Am$ source. The events that induced a charge trigger are shown in red, while those that are under the threshold are shown in greed. The photo-absorption peak can be seen at around 1800 ADC, while the Compton-dominated spectrum can be seen at lower energies. Below $\approx 400\,\mathrm{ADC}$, the individual photo-electron peaks can be clearly observed.
The charge trigger applied here corresponds to approximately 5 photo-electrons, or in the case of this setup 5 keV. The individual photo-electron peaks can be clearly observed up to more than 10 p.e.. In addition, the photo-absorption peak resulting from 59.5 keV absorption in the plastic can be seen around 1800 ADC. The spectrum, as observed in the low gain range also shows this peak, around 180 ADC, indicating that the dynamic range in the low gain reaches up to approximately 1.3 MeV. This is achieved by setting the HG and LG Gain DAC values equal. The range can be optimized for specific purposes by changing these values.

### v2

@@ 2.2.6 Current Trigger Scheme
External triggers can also be provided to the FPGA over an LVDS line. These are used to readout the pedestal or baseline levels of all 64 channels. These events allow to calibrate both the baseline as well as the noise levels as will be discussed later on. For science data, the charge trigger is setup at a higher level such that, in case of 1 channel on the front-end having a significant energy deposition, the FPGA forces peak sensing to commence on all channels and initiates a readout of the ADCs. In addition, the FPGA sends out a signal on the ’trig out’ LVDS line to the back-end. This starts a clock on the back-end during which it waits for data from any other front-end which it combines into overall event packets for the full instrument. To avoid chance coincidence events, for POLAR-2, the clock is set to 100 ns. This allows to detect coincident events within the full detector, while, by using the higher of the two threshold, avoids dark noise induced dead time.
@@
The time threshold is instead setup to record coincident events within a single front-end. The time threshold is set at a level within the dark noise region (typically around 4 or 5 photo-electrons). However, to avoid significant dead time, the ADC readout is only initialized when at least 2 of the time trigger lines are high. This check is performed by the FPGA using the multiplexed time threshold outputs from the 2 Citirocs. This check is performed within 90 nanoseconds. The multiplicity counting by the FPGA also allows to reduce charged particle induced events, as these typically produce events within the detector with a high number of detector channels triggering. In case that within a single front-end the multiplicity is high, the FGPA uses one of the spare LVDS trigger lines to indicate this to the back-end, thereby allowing it to veto any triggers provided by other front-ends.
@@
The trigger logic presented here is optimized to readout a detector with a large number of readout channels in case at least two of these channels exceed a threshold. When used as a spectrometer, this design can be simplified by, for example, only using one of the two thresholds. In addition, it can also be made more complex by, for example requiring at least one channel to be above the charge threshold with a second being required to be above the time threshold. Such a solution can be useful when different types of scintillators are connected to one front-end.
@@ 2.2.7 HVPS
It should be noted that after radiation damage the dark current will significantly increase, see e.g. [28, 24, 29]. For POLAR-2, the yearly radiation dose has been estimated to be 0.0789 Gy/yr (at an altitude of 383 km) through simulations. This would result in an increase in dark current of a factor of 7.2 (for the 6025 or 6050 types would be a factor 15.7 and 15.4, respectively [24]). After one year, the current in one readout system would therefore increase to the order of $1-10mA$, while after several years, when correcting for annealing, this could increase to the $100\,\mathrm{mA}$ level. The LT3482 DC/DC converter can provide up to 200 mA at 70 V over voltage. Therefore, it allows to continue to power the SiPM array even after significant levels of radiation damage.
@@ 3.1 Power Consumption
3.2 Gain Variations
In this design, the signals from the SiPMs arrive at the ASICs where they are amplified. This is done by both a high gain and a low gain amplifier, therefore two amplified signals are produced with different amplitudes. These are finally processed in ADCs which produce the digital values processed in the FPGA. In an ideal detector, one records similar ADC values when two equal input charges are injected in different detector channels. Apart from an equal response of the different channels in the electronics, this also requires an equal sensitivity of the scintillators and SiPMs. While the former can be achieved through careful quality control of the scintillator production and placement, the latter can generally be achieved by optimizing the voltages applied to the various SiPM channels to equalize variations in their gain. For the electronics discussed here, it is important to understand whether an equal charge from the SiPM results in a similar ADC value. For this purpose, charges produced using a pulse generator were injected directly into the electronics just before the ASICs, thereby bypassing the SiPMs and allowing to study the electronics independently. For this test, charges ranging from 20 mV to 360 mV, with incrementing steps of 20 mV were injected in all channels of one ASIC.
The resulting spectrum, containing all the injected charges of the scan, for one channel are shown in the left panel of figure 4. In this figure we can see that each injected charge results in a gaussian distribution of ADC values. The width of this is an indication of the electronic noise which will be discussed in detail in section 3.3. In addition, we see that the lowest charge injection is seen at around 900 ADC, this is a result of a baseline, or pedestal (around 650 ADC for this channel). Finally, it can be observed that at high ADC values the charge induced peaks are closer to each other than at low ADC values. This is an effect of the saturation of the amplifier, an effect which will be discussed and quantified in detail in section 4.2.
Figure 4: Left: The high gain ADC spectrum from channel 0 for a measurement where charges were injected directly into the ASIC. For this test, charges ranging from 20 mV to 360 mV, with incrementing steps of 20 mV were injected. Right: The gain in ADC/mV for the 32 channels of the tested ASIC as measured using direct charge injection.
The distance in ADC between the peaks in the linear regime (below  2000 ADC) can be used to quantify the electronic gain in the system in terms of ADC/mV. The right panel of figure 4 shows the distribution of this gain for the 32 channels of this ASIC. We can see that they vary by about $10\%$. It should be noted, however, that here the electronic gain (the High Gain Amplification from table 1) was equal for all channels. This 8 bit value can be adjusted for each channel to modify the electronic gain by a factor of around 6. Therefore variations of the order of $10\%$ between channels can be removed through optimization if needed.
@@ 3.3 Electronic Noise
To produce scientific data, the system will be triggered to perform a readout of all the ADC values when one, or potentially several, detector channels are above the charge or time threshold. However, in order to measure the baseline as well as the electronic noise, the system can be forced to trigger using the ’trig in’ LVDS line. During standard data taking this can be done, for example, once per second to monitor potential drifts in the baseline due to temperature. Here we use this function to measure the electronic noise. For this purpose approximately 5000 triggers were forced while the high voltage was switched off, thereby only measuring the baseline, and its variations for the 64 channels, both in the high gain and low gain output. An example of the ADC distribution of channel 0, fitted using a Gaussian function can be seen on the left in figure 5. As no HV is applied to the SiPM, the noise seen here is that produced by the electronics on the front-end and injected into the HG line of this ASIC channel. The $\sigma$ values for the 64 channels, for the high gain, can be seen on the right. Similar values were found for the low gain, indicating a typical electronics-induced noise of the order of 5 ADC.
@@
The electronic noise can consist of two components, one which is shared between all the channels in an ASIC, or within the full board, and another which is individual for each channel. The presence of any shared noise component was tested by taking the mean ADC value of the 64 channels and studying its distribution. The resulting distribution was found to have a width consistent with 0 ADC, thereby excluding the presence of a significant shared noise component. The distribution in figure 5 therefore shows the channel individual noise. It should be noted that the same level of noise was found by fitting the distributions in figure 4 acquired through direct charge injection.
@@ 3.4 Rate Capability
The SiPM readout is originally designed to observe transient sources like GRBs. As such it is designed to handle relatively high photon rates. From the POLAR mission, the brightest observed GRB was the short GRB 170127C which, after correcting for dead time, induced a gamma-ray rate of  $\approx\,200\,\mathrm{ph\,cm^{-2}\,s^{-1}}$ for a duration of 100 ms. For the brightest GRB ever observed, 221009A the GRBAlpha collaboration found a flux of $1020\,\mathrm{ph\,cm^{-2}\,s^{-1}}$ in the 50-300 keV range [30]. Using such a rate, and combining it with the predicted effective area of POLAR-2 from [31], we find that, to be capable of observing such transients without a significant amount of dead time, the instrument should be capable of handling trigger rates of hundreds of $\mathrm{triggers/(cm^{2}\,s)}$. Assuming the baseline design, which reads out 64 SiPMs with dimensions of $6\times 6\,\mathrm{mm^{2}}$, this corresponds to being able to handle a trigger rate of the front-end board of the order of several kHz.
The main sources which limit the readout rate of the system are identified as, the ADC readout time, the event packaging time in the FPGA and the readout rate of the board by the back-end. The 14-bit ADC readout is the first bottleneck and puts a lower limit on the possible time between two triggered events. With a clock speed of 1.5 MHz, the total readout time is $9.3\,\mu s$.
@@
The effects can be seen on the left in figure 6 which shows the time between two consecutive events for a data run where thresholds were set such that the system was triggering at high rates on dark noise from the SiPMs. These measurements were taken at room temperature with new SiPMs. Typical dark count rates at the single p.e. level, based on the data sheets of these SiPMs, of several MHz are therefore expected. By setting the thresholds at 3 p.e. trigger rates of several 10’s of kHz are produced. At large times, the peak induced by the back-end at 2 ms can be observed, while at lower times the effects due to the readout system itself. These can be more clearly observed in the inset. The lower limit at $9\,\mu s$ can clearly be observed, while a kink at around $200\,\mu s$, induced by the FGPA can be observed. Fitting the region below the kink using an exponential we can derive a trigger rate of 4 kHz.
@@
To further verify the trigger rate using a controlled trigger rate, charges were directly injected into the setup using the charge injection point on the PCB. The charges, produced by a pulse generator, are injected directly into the ASICs, thereby allowing to test the full electronics chain. When injecting charges at a rate of 1–2 kHz, the correct time between triggers could be recovered. Only when exceeding 5 kHz do we start to see this being distorted. The result of injecting charges at 10 kHz can be observed on the right of figure 6. While for an ideal detector, we would expect to measure 100 $\mu s$ delays, here, for about $60\%$ of the cases, a delay of 200 $\mu s$ is measured, indicating that packets are being lost.
@@ 3.5 Crosstalk
A typical detector using scintillators, SiPMs, and electronics can suffer from various types of crosstalk. Firstly, there can be optical crosstalk between the scintillators. Secondly, crosstalk between microcells in the SiPM occurs. These first two forms of crosstalk depend on the detector design. The third, the electronic crosstalk, depends on the readout design. As the system discussed here is developed to read out 64 independent SiPMs, it is important to avoid any electronic crosstalk between readout lines in the system. A first method to check for electrical cross talk is through the use of injecting charges directly into the PCB. From the charge injection point, the charges go into the ASICs, thereby allowing to check for electrical crosstalk in the ASIC, ADC and in traces towards the FPGA. Even when injecting charges which saturate the LG no signals in adjacent channels were found.
Although positive, this method does not exclude any electrical cross talk from the SiPM towards the ASIC. To verify that no significant crosstalk of this type exists also here, data with large signals from the SiPM can be used. For this purpose, we use data from acquisitions where the system was coupled to an array of GAGG crystals which was illuminated by a 100 keV beam. The correlation between ADC values in different channels was evaluated and studied for crosstalk. To verify that the analysis works, we initially looked at data from two SiPM channels which are geometrically adjacent. The mean ADC value as registered in a geometrically neighboring channel, as a function of the ADC value in the triggering channel, which here was illuminated by the beam, can be seen on the left in figure 7. We see a clear correlation between the two ADC values, indicating that a significant level of crosstalk, either optical or between SiPM channels, exists. This was expected, since the GAGG crystal array is not well optically insulated, thereby resulting in optical crosstalk of $\approx 10\%$.
@@
4 Performance with EJ-248M and MPPC S13361
The SiPM readout system was originally designed as part of the POLAR-2 mission, a large-scale plastic scintillator-based polarimeter [31]. It was therefore initially optimized to read out a densely packed array of plastic scintillator bars. Although originally, EJ-200 was chosen for this purpose, it was found that EJ-248M performed better in terms of light yield despite its lower intrinsic scintillation light yield compared to EJ-200. This was found to be a result of EJ-248M being a relatively harder plastic and therefore allowing a smoother surface quality, resulting in a larger fraction of the light reaching the SiPM. The details of this are discussed in [32], where the details on the optical coupling are also presented along with the calculation of the achieved light yield. In an optimized setup an average of $\approx 1.6$ p.e./keV was measured.
Similarly, the combination of the readout system together with an EJ-248M, as well as an EJ-200 array, was tested during a beam test at the European Synchrotron Radiation Facility (ESRF) of which the details are presented in [31]. While the details on the light yield and spectra can be found in [31], it should be noted that those results were produced using the same electronics, though with a slightly faulty firmware. Here, we present the typical response of the system after a firmware update. For details on the performance of the system with a polarized beam, we refer the reader to [31], while for details on the optical performance for this system, we refer the reader to [32].
Such measurements, as well as those discussed in this and the next section, were performed with an over-voltage of 3.0 V for all 64 channels. To achieve this, the breakdown voltages of all 64 SiPM channels were measured at various temperatures prior to mounting them on the PCB. This over-voltage was selected as it allows for photon detection efficiency of $\approx 50\%$ while keeping the SiPM cross talk below $10\%$  [^4] https://www.hamamatsu.com/content/dam/hamamatsu-photonics/sites/documents/99_SALES_LIBRARY/ssd/s13360_series_kapd1052e.pdf. The POLAR-2 mission relies for its scientific measurements on the detection of coincident energy depositions in 2 detector channels, one of these energy depositions can be as low as several keV. For this mission it is therefore important to keep the energy threshold as low as possible. As this threshold is driven by a combination of the dark noise and the crosstalk, a low crosstalk in the SiPM is more important than a large photon detection efficiency. For a spectrometer with less readout channels, one might consider increasing the over-voltage to prioritize the photon detection efficiency over a higher crosstalk.
@@ 4.1 Spectral Measurements
In order to study the performance of the readout system with an array of EJ-248M scintillators, a full detector, consisting of 64 scintillator bars, was scanned using a collimated ${}^{241}Am$ source. The emission spectrum of this source is dominated by a $59.5\,\mathrm{keV}$ line. The spectrum from one of the detector channels can be seen in figure 8. In this spectrum, which is shown after pedestal subtraction was performed, events that have induced a charge trigger are shown in red, while those that did not are shown in green. It can be seen that the threshold was set so low that a small fraction of events from the dark noise region induced a trigger. This is possible as the threshold has a finite width resulting in a non-zero threshold efficiency at low ADC values. This, in combination with the very high rates in the dark noise region, results in such triggers. Increasing the threshold by approximately 50 ADC fully removes the dark noise-induced triggers.
Figure 8: The energy spectra, in ADC, after pedestal subtraction both in the high gain range (top) and low gain range (bottom). These spectra were obtained when irradiating the readout system connected to a plastic scintillator array at room temperature with an ${}^{241}Am$ source. The events that induced a charge trigger are shown in red, while those that are under the threshold are shown in green. The photo-absorption peak can be seen at around 1800 ADC, while the Compton-dominated spectrum can be seen at lower energies. Below $\approx 400\,\mathrm{ADC}$, the individual photo-electron peaks can be clearly observed.
Within the detector used here, the $59.5\,\mathrm{keV}$ photons are absorbed in a plastic scintillator which, based on its specifications, converts this into approximately 550 optical photons. Based on simulations presented in [32] about 1/3rd of these optical photons reach the SiPM where, as a result of the photon detection efficiency of the SiPM $50\%$ of these are converted to photo-electrons which, the electrical signal from these gets amplified and finally converted into a digital signal shown on the x-axis of figure 8. Overall, based on the various conversions one expects about 1 p.e. per keV deposited. The charge trigger applied here corresponds to approximately 5 p.e., or 5 keV. The individual photo-electron peaks can be clearly observed up to more than 10 p.e.. In addition, the photo-absorption peak resulting from 59.5 keV absorption in the plastic can be seen around 1800 ADC. The spectrum, as observed in the low gain range also shows this peak, around 180 ADC, indicating that the dynamic range in the low gain reaches up to approximately 1.3 MeV. This is achieved by setting the HG and LG Gain DAC values equal. The range can be optimized for specific purposes by changing these values.